# Current affiliation:]Université Paris-Saclay, CNRS, Laboratoire de Physique des Solides, 91405 Orsay, France

### v1

@@ Abstract
We investigated the electronic structure of Co-intercalated 2H-TaS2 using angle-resolved photoemission spectroscopy (ARPES). In the compound Co1/3TaS2, the main electronic bands closely resemble those of pristine 2H-TaS2, with no clear signs of band folding. However, a shallow electron pocket, referred to as the $\beta$-feature, was detected at the Fermi level near the corner of the superlattice Brillouin zone. The surface vs bulk origin of this feature is debated, as it cannot be reproduced using standard DFT calculations. To resolve this, we employed cluster perturbation theory (CPT) to incorporating an exact treatment of strong electron correlations (U) on the cobalt sites, going beyond DFT+U approximation. To further substantiate this, we studied an underdoped sample, Co0.22TaS2, where a reduced charge transfer leads to different Co orbital character near the Fermi level. We find that its electronic structure closely resembles that of undoped 2H-TaS2, and crucially, lacks the $\beta$-feature. Our results demonstrate that the $\beta$-feature is of the bulk origin emerging from the strong electronic correlations where both the Co charge state and long-range crystallographic order play an important role. This work highlights the need for accurate treatment of electron correlations when studying intercalated transition metal dichalcogenides.
@@ I Introduction
Two dimensional magnetic materials are gaining attention because of their potential for spintronic applications and engineering of magnetic heterostructure devices [1]. Among them, transition metal dichalcogenides (TMDs) intercalated by magnetic atoms are extensively being reinvestigated after half a century from their discovery [2, 3]. Their structural diversity, complex and rich electronic and magnetic properties offer a versatile playground to explore intricate interplay among symmetry, magnetic interactions, and electron correlations, that can also be controlled by external fields [3, 4, 5, 6]. Systems intercalated with cobalt attract special attention, mainly due to the large anomalous Hall effect observed in 2H-NbS2 and 2H-TaS2 intercalated compounds [7, 8, 9, 10, 11, 12, 13, 14, 15]. In this paper, we focus on Co1/3TaS2, which features a hexagonal lattice with space group $P6_{3}22$. Intercalated Co atoms occupy octahedral site coordinated by S atoms from layers above and below, and form $\sqrt{3}\times\sqrt{3}$ superlattice rotated by 30 from the $\times1$ unit cell of the parent compound 2H-TaS2 [16, 17, 18, 19]. It is long known that Co1/3TaS2 (also denoted as CoTa3S6) exhibits antiferromagnetic (AF) ordering below $T_{\mathrm{N}}=35$ K [17, 18], though details of its magnetic structure remained elusive [19].
Recently, new progress in the field was achieved when a systematic investigation by Park et al.  [20], revealed two distinct magnetic phases in CoxTaS2 around $x\sim 1/3$. In the overdoped regime ($x\geqslant$0.330) the system hosts AF phase below 35 K, with the magnetic structure found to be coplanar helical order [20]. In contrast, the underdoped regime (0.299 $\leqslant x\leqslant$ 0.325) hosts two distinct AF phases, the first one setting in at $T_{\mathrm{N1}}\approx 38$ K, and second one at $T_{\mathrm{N2}}\approx 26.5$ K accompanied by a large anomalous Hall effect (AHE). Although a small ferromagnetic component is observed below $T_{\mathrm{N2}}$, such weak ferromagnetic moment is considered insufficient to explain the observed large anomalous Hall coefficient [14]. The magnetic structure below $T_{\mathrm{N2}}$ has been uncovered to be a tetrahedral triple-$\mathbf{q}$ magnetic ordering by neutron diffraction experiments [15, 21]. This state arises from the superposition of three coplanar 120 antiferromagnetic modulations, resulting in a non-coplanar spin texture that generates the finite Berry curvature characterized by a net scalar spin chirality and can lead to a spontaneous topological Hall effect (THE).
The substantial changes observed in physical properties of CoxTaS2 with $x\leqslant$ 0.325 compared to $x\geqslant$0.330) cannot be explained by a simple shift of the Fermi level with increasing cobalt concentration. If these differences were purely electronic in origin, they would require a drastic reconstruction of the band structure and Fermi surface topology. Indeed, significant departures from rigid band shift have already been observed in related intercalated compounds such as Co1/3NbS2, Cr1/3NbS2, and Ni1/3NbS2, where angle-resolved photoemission spectroscopy (ARPES) revealed a distinct shallow band near the Fermi level, commonly referred to as the $\beta$ band, likely originating from intercalant Co/Ni orbitals [22, 10, 11, 12, 23]. These findings underscore the critical role of intercalant-derived electronic states in shaping the low-energy electronic structure and their potential impact on topological transport phenomena. In addition to these electronic effects, the observed sensitivity of the magnetic ground state to the Co concentration may also reflect competing magnetic interactions. Notably, the 1/3 intercalation level gives rise to a highly frustrated magnetic environment, where subtle imbalances can tip the system toward different magnetic orders. Slight deviations from this concentration, either by under- or over-doping, appear to lift the frustration and favor one of the competing interactions, leading to qualitatively different magnetic and consequently transport properties.
In this study, we perform ARPES measurements on Co1/3TaS2 and underdoped Co0.22TaS2 to explore how variations in cobalt concentration influence the band dispersion and Fermi surface topology. To complement the experimental data, we also employ theoretical modeling that goes beyond conventional density functional theory (DFT). Specifically, we use cluster perturbation theory (CPT) to account for strong electronic correlations and the role of intercalant-derived states. This extended approach allows us to analyze the origin of the shallow electron band in Co1/3TaS2, commonly referred to as the $\beta$ band, and draw conclusions applicable more generally to magnetically intercalated TMDs.
@@ II Experiment and Methods
Single crystals of Co1/3TaS2 were grown by chemical vapor transport method using iodine as the transport agent [2]. The grown single crystals were characterized by electrical resistivity measurement, X-ray diffraction, energy dispersive X-ray spectroscopy (EDX) and magnetic susceptibility. The EDX analysis indicated 32% intercalation. To achieve this level of intercalation, it was important to put $\sim$ 15% more Co than stoichiometric composition would indicate. Further in text this sample is refered to as Co1/3TaS2. Magnetic susceptibility revealed antiferromagnetic transition temperature at $T_{\mathrm{N}}$ = 37 K. $T_{\mathrm{N2}}$ was not observed in our crystal. The sample does show zero field magnetization below $T_{\mathrm{N}}$ (see Supplemental material) similarly as the double transition samples do below $T_{\mathrm{N2}}$ [14, 21, 15], although the values in our crystals are order of magnitude lower, and anomalous Hall effect is absent. Indeed, contrary to our EDX results, magnetic and transport properties (see Supplemental material) point to slight over doping according to Park et al.  [20]. Temperature dependent electrical resistivity shows a kink at the magnetic transition temperature. Residual resistivity ratio (ratio of the room temperature resistivity to resistivity at 2 K) in our crystals was around 2.2 which is slightly higher than the one reported in double transition samples [14] indicating high sample quality. Analysis of X-ray diffraction data shows that the compound crystallizes in the hexagonal space group No. 182 ($P6_{3}22$), with lattice constants a=5.76Åa=5.76\ $\text{\AA }$ and c=11.99Åc=11.99\ $\text{\AA }$. We have also synthesized a single crystal with 22% Co, as indicated by EDX. For it the starting material contained 33% of Co per Ta. The electrical resistivity exhibited a nonmetallic trend, varying only weakly with temperature. Magnetic susceptibility reveals ferromagnetic like transition around 25 K. The same behavior of 22% intercalate was reported by Liu et al.  [13].
Angle-resolved photoemission spectroscopy (ARPES) measurements were performed at the APE-LE beamline [24] of the Elettra synchrotron. The crystals were oriented along the high-symmetry $\mathrm{\Gamma}$M0 direction using Laue diffraction and then mounted on copper plates with silver epoxy. Clean surfaces were prepared by cleaving inside the ARPES main chamber under ultra-high vacuum (base pressure of $3\times 10^{-10}$ mbar) at 20 K. The incident photon energy ($h\nu$) was varied between 30 and 90 eV. Most measurements were carried out at 20 K, below the magnetic ordering temperature of Co1/3TaS2. To compare the spectra below and above the magnetic transition, additional scans along the $\mathrm{\Gamma}$M0 direction and Fermi surface maps were recorded at 50 K. The photon beam was linearly polarized in the plane of the storage ring and oriented perpendicular to the analyzer slit. Photoelectrons were collected using a hemispherical DA30 analyzer equipped with a multi-channel plate and CCD detector. The overall energy resolution was approximately 30 meV.
@@
The electronic band structure of Co1/3TaS2 was calculated using Density Functional Theory (DFT) as implemented in the Quantum ESPRESSO package [25, 26]. We used a kinetic energy cut-off of 120 Ry for the plane-wave basis set and 1000 Ry for the charge density and potentials. To account for electronic correlations on the cobalt ions, we employed the DFT+U method following the approach of Dudarev et al. [27]. The on-site Hubbard interaction parameter, $U=5.77$ eV, was computed using a linear-response formulation of constrained DFT (LR-cDFT), implemented within the framework of density-functional perturbation theory (DFPT) [28, 29, 30]. The Brillouin zone was sampled using a $\times10$$\times$5 $k$-point mesh without shift. The sharp cutoff of electron occupation at Fermi energy is modeled by a smearing technique proposed by Marzari and Vanderbilt [31], with a broadening width of 0.01 Ry.
@@
Electronic correlations were investigated using Cluster Perturbation Theory (CPT) based on a Wannier-function representation of the spin-polarized DFT+U band structure. The calculations were performed within an energy window of $\pm 1$ eV around the Fermi level. Each cluster consisted of four Wannier nodes representing three Ta and one Co orbitals: the Ta-derived Wannier functions are centered in the triangular plaquettes between Ta atoms, whereas the Co Wannier functions are atom-centered. Local electron interactions on the Co orbitals were included via a Hubbard $U$ term, with double counting corrected by adjusting the on-site energies. The many-body cluster Hamiltonian was solved exactly, while inter-cluster hoppings were treated perturbatively to construct the full CPT Green’s function. The resulting spectral functions were unfolded to the extended Brillouin zone for direct comparison with ARPES measurements.
@@ III Results and discussions
[…]
@@
Fig. 1 shows the ARPES intensity plots of Co1/3TaS2 measured at 20 K with photon energy $h\nu=72$ eV. Panel (a) presents spectra collected along the $\overline{\Gamma}\overline{\mathrm{M}}_{0}$ direction, where $\overline{\mathrm{M}}_{0}$ corresponds to the high-symmetry point in the surface Brillouin zone (BZ) of the host 2H-TaS2, while $\overline{\mathrm{K}}$ marks the high-symmetry point of the Co1/3TaS2 superlattice BZ. The main band features resemble those of pristine 2H-TaS2 [32], with no evident band folding or splitting at the boundaries of the Co1/3TaS2 BZ. Beyond an overall band shift due to charge transfer, slight redistribution of Ta and S atoms upon Co intercalation introduces a modulation into the host lattice. This modulation coupled with some occupational disorder of intercalated Co atoms, results in the observed broadening of the ARPES features compared to the parent compound.
@@
Fig. 1(b) shows the corresponding Fermi surface map. The large black dashed hexagon and the smaller blue dashed hexagon indicate the surface Brillouin zones of 2H-TaS2 and Co1/3TaS2, respectively. As in 2H-TaS2, double walled, cylindrical Fermi surfaces are observed around the $\overline{\mathrm{\Gamma}}$ point ($\alpha_{1}$ and $\alpha_{2}$) and the $\overline{\mathrm{K}}_{0}$ point ($\gamma_{1}$ and $\gamma_{2}$) [33, 34]. Compared to the parent compound, in Co1/3TaS2, the splitting between these double walls is more pronounced around the $\overline{\mathrm{\Gamma}}$ point due to band shifting, and more importantly, Co mediated interlayer hybridization. The same effect was also observed in Co1/3NbS2 [10]. The inner FS ($\alpha_{1}$) is very faint representing a bonding band. Additionally, a notable difference introduced by intercalation is the emergence of a triangular Fermi surface pocket, commonly referred to as the $\beta$-feature", centered at the $\overline{\mathrm{K}}$ point. The band responsible for this Fermi surface pocket is a shallow electron like band around the $\overline{\mathrm{K}}$ point which is marked by the blue dashed line in Fig. 1(a). Similar features have been reported in other intercalated dichalcogenides such as Co1/3NbS2 [10, 11, 12], Ni1/3NbS2 [23], and Cr1/3NbS2 [22, 35].
@@
To study the possible effect of magnetic ordering on the band structure, we also measured the ARPES spectra and the Fermi surface above the magnetic transition ($T_{\mathrm{N}}$ = 37 K) at 50 K. There was no noticeable difference between the spectra at 20 K and 50 K(see Supplemental material). The presence of the $\beta$-feature above and below $T_{\mathrm{N}}$ confirms that the origin of the shallow electron pocket is unrelated to the magnetic transition. The antiferromagnetic TMD series of Co1/3NbS2 and Ni1/3NbS2 also did not exhibit any electronic structure changes across magnetic transitions [36, 23].
@@
Although the $\beta$-feature has been experimentally observed among other transition-metal-intercalated TMDs, its origin remains debated. One possible explanation, as a surface state, has been proposed mainly based on the fact that in Co1/3NbS2 its relative intensity depended on the surface termination or position [11]. However, its variation with excitation photon energy and its observation in soft X-ray ARPES [12] do not convincingly support this interpretation. Furthermore, our DFT calculations considering a slab with two layers of Co1/3TaS2 do not support the surface-effects origin of the $\beta$-feature (see Supplemental material).
@@
Apart from the $\beta$-feature, the DFT calculations successfully reproduce the experimentally observed band structures of Co1/3NbS2 [36] and Ni1/3NbS2 [23]. In particular, the DFT bands unfolded into the large parent-compound BZ match the main band features observed by ARPES. The same is true for Co1/3TaS2 (See Appendix B).
@@
We attribute the origin of the $\beta$-feature to strong correlation effects that cannot be captured by standard DFT+U. In our earlier study of Co1/3NbS2 [36], the feature was addressed within a slave-boson approach where it was shown that strong local Coulomb interactions on the Co sites lead to a renormalization of the Co-host hybridization and a shift of the Co-derived states toward the Fermi level, resulting in the emergence of a narrow, resonance-like, band near the $\overline{\mathrm{K}}$ point, consistent with the experimentally observed $\beta$-feature. The strong correlation origin of this feature was also previously examined within the dynamical mean-field theory (DMFT) framework [37], yielding partial agreement with experiment. While those analyses provided qualitative evidence for the correlation-driven origin of the $\beta$-feature, in the present work, we adopt a more exact treatment of electronic correlations by employing cluster perturbation theory (CPT), initially proposed by Gros et al. [38] and further developed by S n chal et al. [39, 40]. In CPT, the Hubbard interaction is treated exactly at the cluster level, going beyond the mean-field approximation inherent to DFT+U, and capturing momentum-resolved spectral features that are not fully captured by DMFT.
Our CPT calculation is based on a Wannier function description of the band structure obtained from DFT+U for the spin-polarized system. We focus on the energy range around the Fermi level where the $\beta$-feature" occurs, in particular, from 1.0 eV below to 1.5 eV above the Fermi level. In this energy range, seven DFT bands can be accurately fitted (lower graph in Fig. 5) with a tight-binding type Hamiltonian obtained from the Wannier90 package [41] using seven Wannier functions. Several of the most dominant terms of the Wannier Hamiltonian are listed in Table I. While both spin channels share the same band structure, the Wannier representation of DFT results is spin-dependent. The six Wannier functions (spin-independent) are predominantly of Ta $d$-state character and are located in the space forming a triangle between three Ta atoms, not directly at the Ta atom positions (see Fig. 2(a)). The seventh Wannier function is spin-dependent: for electrons with spin up, its location is at the Co atom with spin down, and vice versa. This Wannier function primarily describes the unoccupied band above the Fermi level, while the Ta-derived Wannier functions capture the dispersions of the six bands crossing the Fermi level.

### v2

@@ Abstract
We investigate the electronic structure and Fermi surface of Co1/3TaS2 using angle-resolved photoemission spectroscopy (ARPES) combined with theoretical modeling beyond standard density functional theory (DFT+U). A shallow electron pocket, the so-called $\beta$ feature, is observed at the Fermi level near the corner of the superlattice Brillouin zone, representing the first experimental observation of this feature in an intercalated TaS2 compound. Similar pockets have been reported in $X_{1/3}$NbS2 ($X$ = Co, Cr, Ni), where their surface versus bulk origin remains actively debated. Because conventional DFT+U does not capture this feature, we employ cluster perturbation theory (CPT) to incorporate an explicit treatment of strong electron correlations ($U$) on the Co sites. CPT successfully reproduces the $\beta$ feature, demonstrating its origin from correlation-driven bulk states rather than surface effects. To further substantiate this conclusion, we studied a reduced Co-content sample, Co0.22TaS2, where the reduced charge transfer modifies the Co-derived states near the Fermi level. Its electronic structure remains largely similar to that of pristine 2H-TaS2, showing only a minor overall energy shift and lacking the $\beta$ feature, consistent with disrupted long-range Co ordering and modified orbital character near the Fermi level. We demonstrate that the $\beta$ feature arises from strong local correlations on the Co sites and requires long-range crystallographic order among intercalated Co atoms to maintain coherence. These results highlight the importance of strong electronic correlations in magnetically intercalated transition-metal dichalcogenides and provide a microscopic understanding of features not captured by conventional DFT+U.
@@ I Introduction
Two-dimensional magnetic materials have attracted considerable attention for spintronic applications and the design of magnetic heterostructures [1]. Among them, transition metal dichalcogenides (TMDs) intercalated with magnetic atoms provide a versatile platform, where symmetry, magnetic interactions, and electron correlations intertwine and can be tuned by carrier concentration or external fields [2, 3, 4, 5].
Cobalt-intercalated TMDs have drawn particular interest due to the large anomalous Hall effect (AHE) observed in Co1/3NbS2 and Co1/3TaS2[6, 7, 8, 9, 10, 11, 12, 13, 14]. In Co1/3TaS2, Co atoms occupy octahedral sites coordinated by S atoms and form a $\sqrt{3}\times\sqrt{3}$ superlattice rotated by 30 relative to the $\times1$ unit cell of 2H-TaS2 [15, 16, 17, 18]. The system was long believed to order antiferromagnetically below $T_{\mathrm{N}}\approx 35$ K [16, 17]. However, recent studies have revealed a more complex magnetic phase diagram near $x\approx 1/3$, arising from strong magnetic frustration and a pronounced sensitivity to the Co concentration [19]. For overdoped samples ($x\gtrsim 0.33$), the system exhibits coplanar helical antiferromagnetic order below $T_{\mathrm{N}}\approx 35$ K, whereas in underdoped samples ($x\lesssim 0.325$) two distinct antiferromagnetic transitions occur at $T_{\mathrm{N1}}\approx 38$ K and $T_{\mathrm{N2}}\approx 26.5$ K, including a noncoplanar triple-$\mathbf{q}$ configuration associated with a large AHE and finite scalar spin chirality [14, 20]. Slight deviations from the 1/3 concentration lift the frustration and favor one of the competing interactions, leading to qualitatively different magnetic and consequently transport properties, as well as other characteristics that are crucial for the functionality and application-relevant behavior of these materials.
Such pronounced changes in physical properties cannot be reconciled within a simple rigid-band picture, suggesting an active role of intercalant-derived electronic states in shaping the low-energy electronic structure. Indeed, angle-resolved photoemission spectroscopy (ARPES) studies of related intercalated compounds, Co1/3NbS2, Cr1/3NbS2, and Ni1/3NbS2, have revealed a shallow electron band near the Fermi level, which we refer to here as the $\beta$ band" [21, 9, 10, 11, 22]. Its surface versus bulk origin remains debated, and its microscopic origin is unresolved. Conventional density functional theory (DFT) and its extension DFT+U reproduce the main bands of these intercalated TMDs, but fail to capture the $\beta$ band, indicating that correlation effects beyond the static mean-field description of DFT+U on the intercalant sites play a crucial role.
These observations raise a central question: what is the origin of the shallow $\beta$ band in intercalated TMDs, and how do electronic correlations and intercalant-derived states govern its presence and coherence?
In this work, we combine ARPES measurements on Co1/3TaS2 and underdoped Co0.22TaS2 with theoretical modeling beyond conventional DFT. Using cluster perturbation theory (CPT), we explicitly incorporate local electron correlations on the Co sites and analyze the shallow $\beta$ band in Co1/3TaS2. The underdoped sample allows us to examine how reduced cobalt concentration and the associated disorder in the intercalant lattice affect the coherence of Co-derived bands and modify the low-energy electronic states. These measurements, combined with CPT modeling, provide a framework for understanding how intercalant charge, orbital character, and correlations determine the low-energy electronic structure in magnetically intercalated TMDs.
@@ II Experiment and Methods
Single crystals of Co1/3TaS2 were grown by chemical vapor transport using iodine as the transport agent [23]. The grown crystals were characterized by electrical resistivity, X-ray diffraction, energy dispersive X-ray spectroscopy (EDX), and magnetic susceptibility. EDX analysis indicated 32% intercalation. To achieve this level, it was important to use $\sim$15% more Co than the stoichiometric composition would suggest. In the following, this sample is referred to as Co1/3TaS2. Magnetic susceptibility revealed an antiferromagnetic transition at $T_{\mathrm{N}}=37$ K. $T_{\mathrm{N2}}$ was not observed in our crystal. The sample shows zero-field magnetization below $T_{\mathrm{N}}$ (see Supplemental Material), similarly to double-transition samples below $T_{\mathrm{N2}}$ [13, 20, 14], although the values in our crystals are an order of magnitude lower, and no AHE is observed. Interestingly, despite the EDX results, magnetic and transport measurements (see Supplemental Material) suggest slight overdoping, according to Park et al. [19]. Temperature-dependent electrical resistivity shows a kink at the magnetic transition temperature. The residual resistivity ratio (ratio of room-temperature resistivity to resistivity at 2 K) in our crystals was around 2.2, slightly higher than that reported in double-transition samples [13], indicating high sample quality. Analysis of X-ray diffraction data shows that the compound crystallizes in the hexagonal space group No. 182 ($P6_{3}22$), with lattice constants $a=5.76$ Å and $c=11.99$ Å.
We also synthesized a single crystal with 22% Co, as indicated by EDX. For this sample, the starting material contained 33% Co per Ta. The electrical resistivity exhibited a nonmetallic-like trend, varying only weakly with temperature. Magnetic susceptibility reveals a ferromagnetic-like transition around 25 K. This behavior in the 22% intercalated sample is consistent with previous reports [12].
[…]
@@
The electronic band structure of Co1/3TaS2 was calculated using DFT+U as implemented in the Quantum ESPRESSO package [25, 26]. We used a kinetic energy cut-off of 120 Ry for the plane-wave basis set and 1000 Ry for the charge density and potentials. To account for electronic correlations on the cobalt ions, we employed the DFT+U method following the approach of Dudarev et al. [27]. The on-site Hubbard interaction parameter, $U=5.77$ eV, was computed using a linear-response formulation of constrained DFT (LR-cDFT), implemented within the framework of density-functional perturbation theory (DFPT) [28, 29, 30]. The Brillouin zone was sampled using a $\times10$$\times$5 $k$-point mesh without shift. The sharp cutoff of electron occupation at Fermi energy is modeled by a smearing technique proposed by Marzari and Vanderbilt [31], with a broadening width of 0.01 Ry.
@@
Electronic correlations were investigated using CPT based on a Wannier-function representation of the spin-polarized DFT+U band structure. The calculations were performed within an energy window of $\pm 1$ eV around the Fermi level. Each cluster consisted of four Wannier nodes representing three Ta and one Co orbitals: the Ta-derived Wannier functions are centered in the triangular plaquettes between Ta atoms, whereas the Co Wannier functions are atom-centered. Local electron interactions on the Co orbitals were included via a Hubbard $U$ term, with double counting corrected by adjusting the on-site energies. The many-body cluster Hamiltonian was solved exactly, while inter-cluster hoppings were treated perturbatively to construct the full CPT Green’s function. The resulting spectral functions were unfolded to the extended Brillouin zone for direct comparison with ARPES measurements.
@@ III Results and discussions
Figure 1: (Color online) (a) ARPES intensity plot of Co1/3TaS2 along the $\overline{\Gamma}{\overline{\mathrm{M}}}_{0}$ direction at 20 K measured with $h\nu=72$ eV. The vertical dashed lines indicate the boundaries of the Co1/3TaS2 surface Brillouin zone. (b) Fermi surface map at 20 K measured with $h\nu=72$ eV. The large black hexagon, drawn with a solid line, represents the first Brillouin zone of 2H-TaS2, and the smaller blue hexagon, shown with dashed lines, represents the first Brillouin zone of Co1/3TaS2.
@@
Fig. 1 shows the ARPES intensity plots of Co1/3TaS2 measured at 20 K with $h\nu=72$ eV. Panel (a) presents spectra collected along the $\overline{\Gamma}\overline{\mathrm{M}}_{0}$ direction, where $\overline{\mathrm{M}}_{0}$ corresponds to the high-symmetry point in the surface Brillouin zone (BZ) of the host 2H-TaS2, while $\overline{\mathrm{K}}$ marks the high-symmetry point of the Co1/3TaS2 superlattice BZ. The main band features resemble those of pristine 2H-TaS2 [32], with no obvious band folding or splitting at the boundaries of the Co1/3TaS2 BZ. Beyond an overall band shift due to charge transfer, a slight redistribution of Ta and S atoms upon Co intercalation introduces a modulation in the host lattice. This modulation, coupled with some occupational disorder of intercalated Co atoms, results in the observed broadening of the ARPES features compared to the parent compound.
@@
Fig. 1(b) shows the corresponding Fermi surface map. The large black hexagon and the smaller blue dashed hexagon indicate the surface Brillouin zones of 2H-TaS2 and Co1/3TaS2, respectively. As in 2H-TaS2, double-walled, cylindrical Fermi surfaces are observed around the $\overline{\Gamma}$ point ($\alpha_{1}$ and $\alpha_{2}$) and the $\overline{\mathrm{K}}_{0}$ point ($\gamma_{1}$ and $\gamma_{2}$) [33, 34]. Compared to the parent compound, in Co1/3TaS2, the splitting between these double walls is more pronounced around the $\overline{\Gamma}$ point due to band shifting, and more importantly, Co-mediated interlayer hybridization. The same effect was also observed in Co1/3NbS2 [9]. The inner FS ($\alpha_{1}$) is very faint, representing a bonding band. Additionally, a notable difference introduced by intercalation is the emergence of a triangular Fermi surface pocket, commonly referred to as the $\beta$ feature", centered at the $\overline{\mathrm{K}}$ point. The band responsible for this Fermi surface pocket is a shallow, electron-like band around the $\overline{\mathrm{K}}$ point, which is marked by the blue dashed line in Fig. 1(a). Similar features have been reported in other intercalated dichalcogenides such as Co1/3NbS2 [9, 10, 11], Ni1/3NbS2 [22], and Cr1/3NbS2 [21, 35]. The observation of the $\beta$ feature in Co1/3TaS2 demonstrates that this shallow electron band is not limited to NbS2-based compounds, suggesting a more general, intercalant-driven effect in transition metal dichalcogenides.
@@
To study the possible effect of magnetic ordering on the band structure, we also measured the ARPES spectra and the Fermi surface above the magnetic transition ($T_{\mathrm{N}}=37$ K) at 50 K. There was no observable difference between the spectra at 20 K and 50 K (see Supplemental Material). The presence of the $\beta$ feature above and below $T_{\mathrm{N}}$ confirms that the origin of this shallow electron pocket is independent of the magnetic transition. Similarly, the antiferromagnetic TMDs Co1/3NbS2 and Ni1/3NbS2 also showed no detectable changes in the electronic structure across their magnetic transitions [36, 22].
@@
Although the $\beta$ feature has been experimentally observed in NbS2 intercalates, its origin remains debated. One possible explanation, invoking a surface state, has been proposed primarily because in Co1/3NbS2 its relative intensity varies with surface termination or position [10]. Nevertheless, its dependence on photon energy and observation in soft X-ray ARPES [11] do not convincingly support this scenario. Furthermore, our DFT calculations for a two-layer slab of Co1/3TaS2 do not reproduce the $\beta$ feature as a surface state, indicating that it is unlikely to originate from surface effects (see Supplemental Material).
@@
Apart from the $\beta$ feature, the DFT calculations successfully reproduce the experimentally observed band structures of Co1/3NbS2 [36] and Ni1/3NbS2 [22]. In particular, the DFT bands, when unfolded into the larger parent-compound Brillouin zone, match the main band features observed by ARPES. The same is true for Co1/3TaS2 (see Appendix B).
@@
We attribute the origin of the $\beta$ feature to strong correlation effects that cannot be captured by standard DFT+U. In our earlier study of Co1/3NbS2 [36], the feature was addressed within a slave-boson approach, where it was shown that strong local Coulomb interactions on the Co sites lead to a renormalization of the Co-host hybridization and a shift of the Co-derived states toward the Fermi level, resulting in the emergence of a narrow, resonance-like band near the $\overline{\mathrm{K}}$ point, consistent with the experimentally observed $\beta$ feature. The strong-correlation origin of this feature was also previously examined within the dynamical mean-field theory (DMFT) framework [37], yielding partial agreement with experiment.
While these analyses provided qualitative evidence for the correlation-driven origin of the $\beta$ feature, in the present work we adopt a more exact treatment of electronic correlations by employing CPT, initially proposed by Gros et al. [38] and further developed by S n chal et al. [39, 40]. In CPT, the Hubbard interaction is treated exactly at the cluster level, going beyond the mean-field approximation inherent to DFT+U and capturing momentum-resolved spectral features that are not fully captured by DMFT.
Our CPT calculation is based on a Wannier function description of the band structure obtained from DFT+U for the spin-polarized system. We focus on the energy range around the Fermi level where the $\beta$ feature" appears, in particular, from 1.0 eV below to 1.5 eV above the Fermi level. In this energy range, seven DFT bands can be accurately fitted (for details see lower graph in Fig. 5, shown in Appendix A) with a tight-binding type Hamiltonian obtained from the Wannier90 package [41] using seven Wannier functions. Several of the most dominant terms of the Wannier Hamiltonian are listed in Table I. While both spin channels share the same band structure, the Wannier representation of DFT results is spin-dependent. The six Wannier functions (spin-independent) are predominantly of Ta $d$-state character and are located in the space forming a triangle between three Ta atoms, not directly at the Ta atom positions (see Fig. 2(a)). The seventh Wannier function is spin-dependent: for electrons with spin up, its location is at the Co atom with spin down, and vice versa. This Wannier function primarily describes the unoccupied band above the Fermi level, while the Ta-derived Wannier functions capture the dispersions of the six bands crossing the Fermi level.